### v1

# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

Abstract
In this paper, we consider the convertible code with locally repairable property. We present an improved lower bound on access cost associated with $(r,\delta)$. Then, we provide a general construction of convertible codes with optimal access cost which shows that those codes can be with super-linear length or maximum repairable property. Additionally, employing the known locally repairable codes with super-linear length or maximum repairable property, we provide explicit constructions of convertible codes with super-linear length or maximum repairable property.
Index Terms:
Convertible codes, locally repairable codes, access cost, super-linear length, maximally recoverable codes.
I Introduction
With the rapid expansion of data in large-scale cloud storage and distributed file systems such as Windows Azure Storage and Google File System (GoogleFS), disk failures have become the norm rather than an exception. To against such failures, the simplest approach is to replicate data packets across multiple disks. However, this method incurs significant storage overhead. As an alternative, coding techniques have been developed to enhance storage efficiency.
Considering various aspects of distributed storage system performance, researchers have introduced a range of coding techniques. For instance, to maximize failure tolerance while minimizing redundancy, they have explored maximum distance separable (MDS) codes. To reduce the data transmitted during the repair process, regenerating codes with optimal repair bandwidth have been proposed [10]. To limit the number of nodes involved in data recovery, locally repairable codes have been designed [12]. Over the past decade, significant progress has been made in coding techniques for distributed storage systems based on these performance criteria. For further details, see: Codes with optimal repair bandwidth [10, 14, 45, 49, 44, 27, 25, 43, 50, 29, 17, 16, 8, 46]; Optimal locally repairable codes [18, 12, 37, 41, 1, 22, 28, 48, 5, 6, 7, 15, 39]; Codes with good availability [38, 42, 2, 4, 20]; Codes with optimal update and access properties [43, 35, 26, 9] and references therein.
All these coding schemes are designed under a fixed code rate, which corresponding to a preset failure rate for the storage system. However, device failure rates are inherently dynamic. To solve this problem, we usually preset a low code rate corresponding a worst device failure rate, which may cause waste during the device stateable period. To overcome this drawback, Kadekodi et al. [21] introduced a cost-effective framework that dynamically adapts code rates and redundancy levels to fluctuating device failure rates, thereby optimizing storage efficiency. This process, termed code conversion [32], requires transitioning encoded data from an initial code $[n_{I},k_{I}]$ to a final code $[n_{F},k_{F}]$ with distinct parameters. Code conversion operates in two primary modes:
  1. 1.
    Merge regime: Merging multiple initial codewords into a single final codeword.
  2. 2.
    Split regime: Splitting an initial codeword into multiple final codewords.
The traditional code conversion method involves re-encoding the original data from an initial code $\mathcal{C}_{I}$ to the final code $\mathcal{C}_{F}$. However, this approach requires retrieving a vast number of symbols, decoding the entire dataset, transmitting the data across the network, and subsequently re-encoding it. This results in considerable overhead in terms of network bandwidth, I/O load, and CPU resource utilization within the cluster, thereby impeding the system’s efficiency. To address these limitations, Maturana and Rashmi [32] pioneered the theory of convertible codes enabling parameter transitions without full data reconstruction. Building upon their framework, subsequent studies [33, 34, 31] established fundamental theorems for maximum distance separable (MDS) code conversions. Subsequently, [23] and [11] provided a class of Reed-Solomon (RS) codes with optimal read and write access costs based on a construction of the locally repairable codes (LRCs) [41] and the extended generalized RS codes [19], respectively,
Analogous to MDS codes, storage systems employing LRCs confront the inefficiency of employing code with a fixed rate, but the failure rate of storage devices will change over time. Consequently, the design of LRCs that facilitate efficient code conversions presents both practical and theoretical significance. Maturana and Rashmi [34] first proposed a new construction technique for designing LRCs that can perform code conversion at a lower cost than the default approach by direct computation, projection, piggybacks, projected piggybacks, linear combination, and instance reassignment. Subsequently, Kong [23] established a lower bound on the access cost of code conversion for general LRCs and proposed constructions of LRCs that can perform code conversions with access cost achieving this bound with equality based on LRCs first introduced in [41].
For locally repairable codes, existing results suggest that the code length can grow super-linearly with respect to the alphabet size for some parameter regime, or they allow for recoverable erasure patterns beyond the minimum Hamming distance. In contrast, for convertible codes, known results impose a linear relationship between the alphabet size and the code length while restricting recoverable erasure patterns to sizes smaller than the minimum Hamming distance. This naturally raises the question: Can locally repairable convertible codes achieve super-linear length or support recoverable erasure patterns beyond the minimum Hamming distance? To address this, we first establish an improved lower bound on the access cost associated with $(r,\delta)$ locality. We then present a general construction of convertible codes with optimal access cost, demonstrating that such codes can indeed achieve super-linear length or support recoverable erasure patterns beyond the minimum Hamming distance. Finally, as concrete examples, we construct explicit convertible codes with super-linear length by leveraging known locally repairable codes with the same property. In Table I and Table II, we summarize and compare known related results with our results.
The subsequent sections of the paper are organized as follows: Section II reviews notation and preliminaries. Section III presents improved lower bounds on access cost for locally repairable convertible codes in the merge regime. Section IV proposes a general construction of access-optimal locally repairable convertible codes. Section V proposes explicit access-optimal locally repairable convertible codes with super-linear length or maximally repairable property. Finally, Section VI concludes this paper.
TABLE I: Lower Bounds on The Access Cost for Convertible Codes within Merge Regime ($k_{F}=tk_{I}$)
Code description Read access cost $\rho_{r}$ Write access cost $\rho_{w}$ Reference
$\mathcal{C}_{I}$: $[n_{I},k_{I}]_{q}$ code $\mathcal{C}_{F}$: $[n_{F},k_{F},d]_{q}$-LRC with $(r,\delta)_{a}$-locality $\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\lceil\frac{r\Delta}{r+1}\rceil\right),\mbox{otherwise},\end{cases}$ where $\Delta=n_{F}-2d-(t-1)k_{I}+3-(\lceil\frac{(t-1)k_{I}}{r}\rceil)$ $\rho_{w}\geqslant n_{F}-t(\Delta+d-1)$ [23, Theorem IV.4]
$\mathcal{C}_{I}$: $[n_{I},k_{I}]_{q}$ code $\mathcal{C}_{F}$: $[n_{F},k_{F},d]_{q}$-LRC with $(r,\delta)_{a}$-locality $\rho_{r}\geqslant\begin{cases}k_{F},\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k_{I}+1,\\ t\left(k_{I}-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),\mbox{otherwise},\end{cases}$ where $\Delta=n_{F}-2d-(t-1)k_{I}+2-(\lceil\frac{(t-1)k_{I}}{r}\rceil-1)(\delta-1)$ $\rho_{w}\geqslant n_{F}-t(\Delta+d-1)$ Theorem 2
TABLE II: Known Constructions of $(t,1)_{q}$ LRCCs within Merge Regime with Parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F}=tk_{I},r_{F},\delta_{F})_{a}$
Construction Restrictions Field size $q$ requirement Additional properties
Access optimal LRCCs [23, Construction III,IV] $r_{I}=r_{F}=r$ $\delta_{I}=\delta_{F}=2$ $r|k_{I},r|k_{F}$ $(r+1)|n_{I},(r+1)|n_{F}$ $\frac{k_{I}(r+1)}{r}~{}|~{}(q-1)$ and $q\geqslant\frac{k_{I}(r+1)}{r}\max\{t+1,\lceil\frac{rn_{I}}{(r+1)k_{I}}\rceil+1\}+1$
General access optimal LRCCs (Theorem 3) $t|\ell,\ell/t>g-\ell\geqslant h\geqslant 0$ $r_{I}=r_{F}=r$ $\delta_{I}=\delta_{F}=\delta\geqslant 2$ $n_{I}=(\ell/t+g-\ell)(r+\delta-1)$ $n_{F}=(\ell+h)(r+\delta-1)$ $k_{I}=\ell r/t,k_{F}=\ell r$ Determined by the field size requirement of $\mathcal{C}_{F}$ Examples: super-linear length (Corollary 2) maximum recoverable (Corollary 4)
LRCs with information locality [47] $\mathcal{C}_{I}$: $[20,12]$-LRC with $(2,2)$-locality $\mathcal{C}_{F}$: $[16,12]$-LRC with $(6,2)$-locality, or vice versa (In this box, we show the access cost.) $6$(read) $+2$(write), $\mathcal{C}_{I}\to\mathcal{C}_{F}$, $10$(read) $+6$(write), $\mathcal{C}_{F}\to\mathcal{C}_{I}$. changed locality
II Preliminary
We first introduce some notation used throughout this paper.
  • For a positive integer $a$, let $[a]$ be the set $\{1,2,\dots,a\}$;
  • For a prime power $q$, let $\mathbb{F}_{q}$ be the finite field with $q$ elements and $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$;
  • For two vector spaces $V_{1},V_{2}$ over $\mathbb{F}_{q}$, denote $V_{1}\otimes V_{2}=\{(v_{1},v_{2}):v_{1}\in V_{1},v_{2}\in V_{2}\}$;
  • For a finite set $S$, let $|S|$ denote its cardinality;
  • An $[n,k,d]_{q}$ linear code $\mathcal{C}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with the minimal Hamming distance $d$;
  • For a set $S\subseteq[n]$, let $\mathcal{C}|_{S}$ denote the punctured code formed by all the codewords of $\mathcal{C}$ that delete symbols $c_{j}$ for all $j\in[n]\setminus S$;
  • Let $O_{m\times n}$ and $I_{m\times n}$ denote the $m\times n$ zero matrix and identity matrix, respectively. If $m,n$ are implicit from context, we simply write $O$ and $I$ for short;
  • For a matrix
    $H=\begin{pmatrix}A_{1}&O&\cdots&O\\ O&A_{2}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&A_{g}\\ B_{1}&B_{2}&\cdots&B_{g}\end{pmatrix},$ (1)
    let $H||_{\{i_{1},i_{2},\dots,i_{t}\}}$ denote the submatrix of $H$ as
    $H||_{\{i_{1},i_{2},\dots,i_{t}\}}\triangleq\begin{pmatrix}A_{i_{1}}&O&\cdots&O\\ O&A_{i_{2}}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&A_{i_{t}}\\ B_{i_{1}}&B_{i_{2}}&\cdots&B_{i_{t}}\end{pmatrix},$ (2)
    where $i_{1}<i_{2}<\dots<i_{t}$ and $i_{1},i_{2},\dots,i_{t}\in[g]$;
  • Let $\top$ be the transpose operator.
It is well known that the following lemma characterizes the distance property.
Lemma 1 ([19, Theorem 1.5.7]):
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code. For any set $S\subseteq[n]$, if $n-|S|<d$, then the punctured code $\mathcal{C}|_{S}$ has dimension $k$.
II-A Locally Repairable Codes
Definition 1 ([12],[36]):
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code. The $i$th code symbol $c_{i}$ of $\mathcal{C}$ is called to have $(r,\delta)$-locality if there exists a subset $S_{i}\subset[n]$ (an $(r,\delta)$-repair set) satisfying
  • $i\in S_{i}$ and $|S_{i}|\leqslant r+\delta-1$,
  • the minimum distance of the code $\mathcal{C}|_{S_{i}}$ is at least $\delta$.
For an $[n,k,d]_{q}$ linear code $\mathcal{C}$, if all the $n$ code symbols have $(r,\delta)$-locality, then it is said to have all symbol $(r,\delta)$-locality, denoted as locally repair codes (LRCs) with $(r,\delta)_{a}$-locality. For a LRC with $(r,\delta)_{a}$-locality, Lemma 1 and Definition 1 imply that there exists a subset $\mathrm{Loc}(i)\subseteq S_{i}$ with $|\mathrm{Loc}(i)|\leqslant r$ and $i\not\in\mathrm{Loc}(i)$ such that
$c_{i}=\sum_{j\in\mathrm{Loc}(i)}\lambda_{j}c_{j},$
where $\lambda_{j}\in\mathbb{F}_{q}^{*}$ for all $j\in\mathrm{Loc}(i)$. Note that, for given $i\in[n]$, $\mathrm{Loc}(i)$ may be not unique. In the following, we use the notation $\mathrm{Loc}(i)$ to denote one of the sets with the above property for the $i$th code symbol $c_{i}$.
An upper bound on the minimum Hamming distance of linear codes with $(r,\delta)_{a}$-locality was derived as follows.
Lemma 2 ([12],[36]):
The minimum distance of an $[n,k,d]_{q}$ code $\mathcal{C}$ with $(r,\delta)_{a}$-locality is upper bounded by
$d\leqslant n-k+1-\left(\left\lceil\frac{k}{r}\right\rceil-1\right)(\delta-1).$ (3)
An LRC code is called optimal if the upper bound (3) is achieved. The repair sets of linear code $\mathcal{C}$ with $(r,\delta)_{a}$-locality has the following important property.
Lemma 3 ([40, Theorem 9]):
Let $\mathcal{C}$ be an optimal linear LRC with $(r,\delta)_{a}$-locality. If $r\mid k$ and $r<k$, then there are $t$ repair sets $S_{i_{1}},\dots,S_{i_{t}}$ are mutually disjoint where $S_{i_{j}}\subseteq[n]$ for $j\in[t]$, and $|S_{i_{j}}|=r+\delta-1$ for all $j\in[t]$, and the punctured code $\mathcal{C}|_{S_{i_{j}}}$ is an $[r+\delta-1,r,\delta]$ MDS code. In particular, we have $(r+\delta-1)\mid n$, and $t=\frac{n}{r+\delta-1}$.
II-B Convertible Codes
Recently, [11] generalizes the framework for studying code conversions proposed by [32], allowing all initial and final codes to have arbitrary parameters and specifying the conversion procedure.
Definition 2 (Generalized Convertible Code [11]):
Let $t_{1},t_{2}$ be two positive integers. A $(t_{1},t_{2})_{q}$ generalized convertible code $\bm{\mathcal{C}}$ over $\mathbb{F}_{q}$ consists of:
1) $t_{1}$ initial codes $\mathcal{C}_{I_{1}},\mathcal{C}_{I_{2}},\cdots,\mathcal{C}_{I_{t_{1}}}$ and $t_{2}$ final codes $\mathcal{C}_{F_{1}},\mathcal{C}_{F_{2}},\cdots,\mathcal{C}_{F_{t_{2}}}$, where $\mathcal{C}_{I_{i}}$ is an $[n_{I_{i}},k_{I_{i}}]_{q}$ code, and $\mathcal{C}_{F_{j}}$ is an $[n_{F_{j}},k_{F_{j}}]_{q}$ code, such that
$\sum_{i\in[t_{1}]}k_{I_{i}}=\sum_{j\in[t_{2}]}k_{F_{j}};$
2) A conversion procedure is determined by a function $\sigma=(\sigma_{1},\sigma_{2},\cdots,\sigma_{t_{2}})$ from $\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}\rightarrow\bigotimes_{j\in[t_{2}]}\mathcal{C}_{F_{j}}$, i.e.,
[…]
Specifically, a concrete conversion procedure is realized in three steps.
Step 1: Choosing unchanged symbols from initial codes. For $j\in[t_{2}]$, the conversion procedure keeps some code symbols from each initial code $\mathcal{C}_{I_{i}},1\leqslant i\leqslant t_{1}$. Denote the indices of these symbols as $\mathcal{U}_{1,j},\mathcal{U}_{2,j},\cdots,\mathcal{U}_{t_{1},j}$, where $\mathcal{U}_{i,j}\subseteq\{i\}\times[n_{I_{i}}]$ satisfies
$\mathcal{U}_{i,\tau_{1}}\cap\mathcal{U}_{i,\tau_{2}}=\emptyset\,\,\text{ for }\tau_{1}\neq\tau_{2}\in[t_{2}].$
For the convenience of analysis, we use two-dimensional indices to denote the code symbols, i.e., the codeword $,\bm{c}_{I_{i}}=(c_{i,1},c_{i,2},$$\cdots,c_{i,n_{I_{i}}})\in\mathcal{C}_{I_{i}}$ for $i\in[t_{1}]$.
Step 2: Reading symbols from initial codes. For $j\in[t_{2}]$, the function $\sigma_{j}$ reads some code symbols from each initial code $\mathcal{C}_{I_{i}},1\leqslant i\leqslant t_{1}$, whose indices are denoted by $\mathcal{R}_{1,j},\mathcal{R}_{2,j},\cdots,\mathcal{R}_{t_{1},j}$ where $\mathcal{R}_{i,j}\subseteq\{i\}\times[n_{I_{i}}]$, respectively.
Step 3: Conversion. For any $j\in[t_{2}]$, choose $u_{j}$ unchanged symbols $\mathcal{C}_{I_{1}}|_{\mathcal{U}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{U}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{U}_{t_{1},j}}$. And then apply $\sigma_{j}$ to the read symbols $\mathcal{C}_{I_{1}}|_{\mathcal{R}_{1,j}},\mathcal{C}_{I_{2}}|_{\mathcal{R}_{2,j}},\cdots,\mathcal{C}_{I_{t_{1}}}|_{\mathcal{R}_{t_{1},j}}$ to generate $n_{F_{j}}-u_{j}$ symbols
$\mathcal{C}_{F_{j}}|_{\mathcal{W}_{j}}=\left\{\sigma_{j}(\bm{x}):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}}|_{\mathcal{R}_{i,j}}\right\},$
which are called written symbols, where $\mathcal{W}_{j}\triangleq\{w_{j}\}\times[n_{F_{j}}-u_{j}]$ and $w_{j}\triangleq t_{1}+j$.
Finally, the final code $\mathcal{C}_{F_{j}}$ is given by
$\mathcal{C}_{F_{j}}$ =$=\left\{(\bm{x}_{1},\sigma_{j}(\bm{x}_{2})):\bm{x}\in\bigotimes_{i\in[t_{1}]}\mathcal{C}_{I_{i}},\bm{x}_{1}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}},\bm{x}_{2}=\bm{x}|_{\bigcup_{i\in[t_{1}]}\mathcal{R}_{i,j}}\right\}$
for each $1\leqslant j\leqslant t_{2}$, where we assume the code symbols of $\mathcal{C}_{F_{j}}$ are indexed by the set
$\mathcal{F}=\bigcup_{i\in[t_{1}]}\mathcal{U}_{i,j}\cup\mathcal{W}_{j}$ (4)
such that for every $t_{1}$ initial codewords $\bm{c}_{I_{i}}\in\mathcal{C}_{I_{i}}$, $i\in[t_{1}]$, there are $t_{2}$ unique final codewords $\bm{c}_{F_{j}}\in\mathcal{C}_{F_{j}}$ for $j\in[t_{2}]$ satisfying
$\bm{c}_{I_{i}}|_{\mathcal{U}_{i,j}}=\bm{c}_{F_{j}}|_{\mathcal{U}_{i,j}}.$
When initial codes and final codes are both MDS codes, the corresponding convertible code $\bm{\mathcal{C}}$ is called an MDS convertible code. When initial codes and final codes are both LRCs, the corresponding convertible code $\bm{\mathcal{C}}$ is called an locally repairable convertible code (LRCC). In this paper, we mainly consider that all initial codes are with same parameters in the merge regime, i.e., code length $n_{I}$, dimension $k_{I}$ and $(r_{I},\delta_{I})$-locality, which is called a $(t,1)_{q}$ LRCC with parameters $(n_{I},k_{I},r_{I},\delta_{I};n_{F},k_{F},r_{F},\delta_{F})_{a}$ where the subscript $a$ means the initial and final codes are with all symbols locality. If some parameters are unknown, i.e., $r_{I},\delta_{I}$, then we say the above LRCC is with parameters $(n_{I},k_{I},*,*;n_{F},k_{F},r_{F},\delta_{F})_{a}$.
One indicator of the quality of a convertible code is the read and write access costs, which measure the number of symbols affected by the conversion procedure.
Definition 3 (Access Cost [32][11]):
For a $(t_{1},t_{2})_{q}$ generalized convertible code $\bm{\mathcal{C}}$, the access cost $\rho$ is defined as the sum of read access cost $\rho_{r}=\sum_{i\in[t_{1}]}|\bigcup_{j\in[t_{2}]}\mathcal{R}_{i,j}|$ and write access cost $\rho_{w}=\sum_{j\in[t_{2}]}(n_{F_{j}}-\sum_{i\in[t_{1}]}|\mathcal{U}_{i,j}|)$, where the read access cost is the number of symbols in initial codes used in the conversion procedure functions and the write access cost counts the number of symbols written in the final codes.
In [23], Kong proposed a lower bound on the access cost of a LRCC with the condition that the final code $\mathcal{C}_{F}$ is an $[n,k,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
Theorem 1 ([23, Theorem IV.4]):
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ convertible code with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. Assume that $\mathcal{C}_{F}$ is a linear code with minimum distance $d$. Then, the write access cost is
$\rho_{w}\geqslant t\left(d+(t-1)k+\left\lceil\frac{(t-1)k}{r}\right\rceil-2\right)-(t-1)n_{F}.$
Write $\Delta=n_{F}-2d-((t-1)k+\lceil\frac{(t-1)k}{r}\rceil)+3$. Then, the read access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\text{ or }d>n_{I}-k+1;\\ t(k-\lceil\frac{r\Delta}{r+1}\rceil),&\mbox{otherwise}.\end{cases}$
Remark 1:
When $r=k$, the preceding bound is the same as the lower bound of access cost for MDS convertible codes proposed by [32, Theorem 9], since the $[n,k,d]_{q}$ MDS codes are optimal LRCs with $(k,d)_{a}$-locality.
III Improved Lower Bound on Access Cost
In this section, we improve Theorem 1 by establishing a new lower bound that depends on both $r$ and $\delta$. We first derive an upper bound on the number of unchanged symbols for each initial code.
Lemma 4:
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. Assume that the final code $\mathcal{C}_{F}$ has minimum distance $d$. Then, for every $1\leqslant i\leqslant t$, we have
$|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1).$
Proof:
Without loss of generality, assume $i=1$. Let $\tilde{\mathcal{C}}$ be the subcode of the final code $\mathcal{C}_{F}$ that is obtained by converting the initial codewords $\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{t}$, where $\bm{c}_{1}=\bm{\alpha}$ is a constant vector and $\bm{c}_{i}\in\mathcal{C}_{I_{i}}$ for $i\in[t]\setminus\{1\}$. Let $d(\tilde{\mathcal{C}})$ denote the minimum Hamming distance of $\tilde{\mathcal{C}}$. Then we have $d(\tilde{\mathcal{C}})\geqslant d$ and $|\tilde{\mathcal{C}}|=q^{(t-1)k}$. Define
$N\triangleq\{i\in\mathcal{F}:i\in\mathcal{U}_{1,1}\text{ or exists a set }\mathrm{Loc}(i)\subseteq\mathcal{U}_{1,1}\},$
where $\mathcal{F}$ is defined by (4) and $N^{C}\triangleq\mathcal{F}\setminus N$. Then, for each $i\in N^{C}$ and corresponding $\mathrm{Loc}(i)$, we claim that $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$. Otherwise, assume that there is a $j\in N^{C}$ such that $\mathrm{Loc}(j)\subseteq N$. Then, each symbol whose index is in $\mathrm{Loc}(j)$ can be recovered by symbols with indices in $\mathcal{U}_{1,1}$, implying that the $j$th symbol can be recovered by symbols with indices in $\mathcal{U}_{1,1}$. This leads to $j\in N$, a contradiction.
Let $\bm{v}=\sigma_{1}(\bm{\alpha},\bm{c}_{2},\dots,\bm{c}_{t})\in\tilde{\mathcal{C}}$ be the codeword converted from codewords $\bm{\alpha},\bm{c}_{2},\dots,\bm{c}_{t}$, where $\sigma_{1}$ is the conversion procedure defined in Definition 2. Note that $\bm{v}|_{\mathcal{U}_{1,1}}=\bm{\alpha}|_{\mathcal{U}_{1,1}}$, and symbols in $\bm{v}|_{N}$ can be recovered by symbols in $\bm{v}|_{\mathcal{U}_{1,1}}$. Thus, $\bm{v}|_{N}$ is fully determined by $\bm{\alpha}|_{\mathcal{U}_{1,1}}$. This implies that $\tilde{\mathcal{C}}|_{N}$ is a constant vector determined by $\bm{\alpha}|_{\mathcal{U}_{1,1}}$. Therefore, we have $d(\tilde{\mathcal{C}})=d(\tilde{\mathcal{C}}|_{N^{C}})$ and $|\tilde{\mathcal{C}}|=|\tilde{\mathcal{C}}|_{N^{C}}|=q^{(t-1)k}$.
On the other hand, since for each $i\in N^{C}$, $\mathrm{Loc}(i)$ satisfies $\mathrm{Loc}(i)\cap N^{C}\neq\emptyset$, and $\mathrm{Loc}(i)\subseteq S_{i}$, where $S_{i}$ is an $(r,\delta)$-repair set for the code symbol $c_{i}$, we have $d(\tilde{\mathcal{C}}|_{S_{i}})\geqslant d(\mathcal{C}_{F}|_{S_{i}})\geqslant\delta$ and $\tilde{\mathcal{C}}|_{S_{i}\cap N}$ is a constant vector. It follows that
$d(\tilde{\mathcal{C}}|_{N^{C}\cap S_{i}})=d(\tilde{\mathcal{C}}|_{S_{i}})\geqslant\delta.$
Hence, $\tilde{\mathcal{C}}|_{N^{C}}$ can be regarded as an $(n_{F}-|N|,(t-1)k)$-LRC with $(r,\delta)_{a}$-locality. By Lemma 2, we obtain
$\begin{split}d(\tilde{\mathcal{C}})&=d(\tilde{\mathcal{C}}|_{N^{C}})\\ &\leqslant(n_{F}-|N|)-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1).\end{split}$ (5)
The desirable result follows from $d(\tilde{\mathcal{C}})\geqslant d$, $|\mathcal{U}_{1,1}|\leqslant|N|$ and (5). ∎
Next, we provide a lower bound for the number of read symbols from each initial code.
Lemma 5:
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$. For each $1\leqslant i\leqslant t$, define
$\Delta_{i}\triangleq|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1.$ (6)
Let $B$ be the largest subset of $\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}$ such that
  • (i)
    $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ with $B_{t_{1}}\cap B_{t_{2}}=\emptyset$, $t_{1}\neq t_{2}$ and $\tau\leqslant\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$;
  • (ii)
    For each $1\leqslant t\leqslant\tau$ and $j_{1},j_{2}\in B_{t}$ with $j_{1}\neq j_{2}$, we may set $\mathrm{Loc}(j_{1})$ and $\mathrm{Loc}(j_{2})$ as the same set, i.e., $\mathrm{Loc}(j_{1})=\mathrm{Loc}(j_{2});$
  • (iii)
    For each $1\leqslant t\leqslant\tau$ and $j\in B_{t}$, we have one possible $\mathrm{Loc}(j)$ satisfying
    $\mathrm{Loc}(j)\cap B=\emptyset.$
Then, for each $1\leqslant i\leqslant t$, we have
$|\mathcal{R}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta_{i}\leqslant 0;\\ k-\Delta_{i}+|B|,&\mbox{otherwise}.\end{cases}$
Moreover, if $d>n_{I}-k+1$, then $\Delta_{i}\leqslant 0$ for $1\leqslant i\leqslant t$ and therefore, $|\mathcal{R}_{i,1}|\geqslant k$.
Proof:
Without loss of generality, assume $i=1$. Let $\tilde{\mathcal{C}}$ be the subcode of the final code $\mathcal{C}_{F}$ that arises from converting the initial codewords $\bm{c}_{1},\bm{c}_{2},\dots,\bm{c}_{t}$, where $\bm{c}_{i}=\bm{\beta}_{i}$ are constant vectors for $i=2,\dots,t$, and $\bm{c}_{i}\in\mathcal{C}_{I_{i}}$ for $i\in[t]$. Let $d(\tilde{\mathcal{C}})$ denote the minimum distance of $\tilde{\mathcal{C}}$. Then we have $d(\tilde{\mathcal{C}})\geqslant d$ and $|\tilde{\mathcal{C}}|=q^{k}$.
For any $\bm{v}\in\tilde{\mathcal{C}}$, each symbol of $\bm{v}$ is either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}},\bm{\beta}_{2}|_{\mathcal{U}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{U}_{t,1}}$ or determined by symbols from $\bm{c}_{1}|_{\mathcal{R}_{1,1}},\bm{\beta}_{2}|_{\mathcal{R}_{2,1}},\dots,\bm{\beta}_{t}|_{\mathcal{R}_{t,1}}$. Therefore, the non-constant symbols of the codewords in $\tilde{\mathcal{C}}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$ or determined as functions of symbols in $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$.
According to Lemma 1, we will pick a subset $S\subseteq\mathcal{F}$ of size at least $n_{F}-d+1$ and prove the lower bound by analyzing the relationships among codeword symbols in $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$, $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$ and $\bm{v}|_{S}$.
Case 1: $\Delta_{1}\leqslant 0$, where $\Delta_{1}$ is defined by (6). In this case, $\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ has size at least $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$ by $|\mathcal{F}|=n_{F}$. Take $S=\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$, we have $|\tilde{\mathcal{C}}|_{S}|=|\tilde{\mathcal{C}}|=q^{k}$. Note that $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}=\bm{v}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$. Thus, by $S\cap(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})=\emptyset$, non-constant symbols in $\bm{v}|_{S}$ are functions of symbols in $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$. Therefore, $|\tilde{\mathcal{C}}|_{S}|$ is upper bounded by $q^{|\mathcal{R}_{1,1}|}$, which implies that $|\mathcal{R}_{1,1}|\geqslant k$.
Case 2: $\Delta_{1}>0$. Let $N=\bigcup_{j\in B}\mathrm{Loc}(j)$. Since the final code $\mathcal{C}_{F}$ has $(r,\delta)_{a}$-locality, we have
$|B\cup N|\leqslant\left\lfloor\frac{\Delta_{1}}{r+\delta-1}\right\rfloor\cdot(r+\delta-1)\leqslant\Delta_{1},$
where the first inequality holds by the conditions (i) and (ii). Note that
$|(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\setminus(B\cup N)|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}|-|B\cup N|\geqslant|\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}|-\Delta_{1}=d-1.$
Therefore, we can obtain an $(n_{F}-d+1)$-subset $S\subseteq\mathcal{F}$ by removing any $d-1$ elements in $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\setminus(B\cup N)$ from $\mathcal{F}$. Note that both $B\cup N$ and $\mathcal{F}\setminus(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ are subsets of $S$. Thus, by $n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}$.
Let $S_{1}=S\cap(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$. Then, $|S_{1}|=|(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})|-d+1=\Delta_{1}$. Since $B\subseteq(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})$ and $B\cup N\subseteq S$, we have $B\subseteq S_{1}$. Moreover, by $N\cap B=\emptyset$ from the condition (iii), we have $N\subseteq S\setminus B$. Thus, $\bm{v}|_{B}$ can be recovered from the symbols $\bm{v}_{N}$ and then from the symbols in $\bm{v}_{S\setminus B}$.
Recall that $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}=\bm{v}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$. Thus, by $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\cap S=S_{1}$, the symbols in $\bm{v}|_{S}$ that are inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}}$ originate from $\bm{c}_{1}|_{S_{1}}$. This implies that the non-constant symbols in $\bm{v}|_{S\setminus B}$ are determined by $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$ and $\bm{c}_{1}|_{S_{1}\setminus B}$. Therefore, $|\tilde{\mathcal{C}}|_{S}|$ is upper bounded by
$q^{|\mathcal{R}_{1,1}\cup(S_{1}\setminus B)|}=q^{|\mathcal{R}_{1,1}|+|S_{1}|-|B|},$
where the equality follows from $S_{1}\setminus B\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$, $(\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1})\cap\mathcal{R}_{1,1}=\emptyset$, and $B\subseteq S_{1}$. This leads to $|\mathcal{R}_{1,1}|+|S_{1}|-|B|\geqslant k$, which further implies that $|\mathcal{R}_{1,1}|\geqslant k-\Delta_{1}+|B|$.
For the second part, we consider the case such that $d>n_{I}-k+1$. In this case, we claim that one must have $\Delta_{1}\leqslant 0$. Otherwise, assume that $\Delta_{1}>0$ and let $D\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$ such that $|D|=d-1$. Let $S=\mathcal{F}\setminus D$. By $|S|=n_{F}-d+1\geqslant n_{F}-d(\tilde{\mathcal{C}})+1$, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}$. Note that for each $\bm{v}\in\tilde{\mathcal{C}}$, the non-constant symbols of $\bm{v}|_{S}$ are either inherited from $\bm{c}_{1}|_{\mathcal{U}_{1,1}\setminus D}$ or determined by $\bm{c}_{1}|_{\mathcal{R}_{1,1}}$. Therefore, we have $|\tilde{\mathcal{C}}|_{S}|=q^{k}\leqslant q^{|(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}|}$. Since $D\subseteq\mathcal{U}_{1,1}\setminus\mathcal{R}_{1,1}$, we have $(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}=(\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1})\setminus D$ and $|(\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1})\setminus D|=|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|$. This leads to
$|(\mathcal{U}_{1,1}\setminus D)\cup\mathcal{R}_{1,1}|=|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|\geqslant k.$
On the other hand, since $\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}\subseteq\{1\}\times[n_{I}]$, we have
$k\leqslant|\mathcal{U}_{1,1}\cup\mathcal{R}_{1,1}|-|D|\leqslant n_{I}-d+1.$
This implies that $d\leqslant n_{I}-k+1$, which is a contradiction. Thus, when $d>n_{I}-k+1$, we always have $\Delta_{1}\leqslant 0$ and therefore, $|\mathcal{R}_{1,1}|\geqslant k$, which completes the proof. ∎
Remark 2:
For $\delta=2$, the condition (ii) can be satisfied only under the case $|B_{i}|=1$ for $1\leqslant i\leqslant\tau$.
The lower bound in Lemma 5 is affected by the size of the subset $|B|$. In the following, we show that there always exists a set $B$ with large cardinality.
Lemma 6:
Assume that $\mathcal{C}_{F}$ is a linear $[n_{F},tk]_{q}$ LRC with $(r,\delta)_{a}$-locality. For any $\tilde{B}\subseteq[n_{F}]$, there is a subset $B\subseteq\tilde{B}$ of size at least $(\delta-1)\lfloor\frac{|\tilde{B}|}{r+\delta-1}\rfloor$ such that
  • (i)
    $B=\cup_{1\leqslant t\leqslant\tau}B_{t}$ with $B_{t_{1}}\cap B_{t_{2}}=\emptyset$, $t_{1}\neq t_{2}$ and $\tau\leqslant\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$;
  • (ii)
    For each $1\leqslant t\leqslant\tau$ and $j_{1},j_{2}\in B_{t}$ with $j_{1}\neq j_{2}$, we may set $\mathrm{Loc}(j_{1})$ and $\mathrm{Loc}(j_{2})$ as the same set, i.e., $\mathrm{Loc}(j_{1})=\mathrm{Loc}(j_{2});$
  • (iii)
    For each $1\leqslant t\leqslant\tau$ and $j\in B_{t}$, we have one possible $\mathrm{Loc}(j)$ satisfying
    $\mathrm{Loc}(j)\cap B=\emptyset.$
Proof:
We may find a set $B$ by the following process. For $\tau=1$ and $i_{1}\in\tilde{B}$, let $\tilde{B}_{i_{1}}=\tilde{B}\cap S_{i_{1}}$,
$B_{i_{1}}=\begin{cases}B^{*}\subseteq\tilde{B}_{i_{1}},&\text{ if }|\tilde{B}_{i_{1}}|\geqslant\delta-1\text{ and }|B^{*}|=\delta-1,\\ \tilde{B}_{i_{1}},&\text{ if }|\tilde{B}_{i_{1}}|<\delta-1.\end{cases}$ (7)
and set
$\tilde{B}\leftarrow\tilde{B}\setminus S_{i_{1}}.$ (8)
For $1<\tau\leqslant|\tilde{B}|$, if $|\tilde{B}|\neq\emptyset$ then we set $\tilde{B}_{i_{\tau}}=\tilde{B}\cap S_{i_{\tau}}$ for a $i_{\tau}\in\tilde{B}$
$B_{i_{\tau}}=\begin{cases}B^{*}\subseteq\tilde{B}_{i_{\tau}},&\text{ if }|\tilde{B}_{i_{\tau}}|\geqslant\delta-1\text{ and }|B^{*}|=\delta-1,\\ \tilde{B}_{i_{\tau}},&\text{ if }|\tilde{B}_{i_{\tau}}|<\delta-1,\end{cases}$ (9)
and
$\tilde{B}\leftarrow\tilde{B}\setminus S_{i_{\tau}}.$ (10)
For $1<\tau\leqslant|\tilde{B}|$, if $\tilde{B}=\emptyset$, then set
$B=\bigcup_{1\leqslant j\leqslant\tau-1}B_{i_{j}}.$ (11)
By (8) and (10), we have $B$ satisfies the condition (i). By (7) and (9), we have $B$ fulfills the condition (ii) since we can find a $\textrm{Loc}(\ell)\subseteq S_{i_{j}}\setminus B_{i_{j}}$ such that the code symbol $c_{\ell^{\prime}}$ can be recovered by the code symbols with index $\textrm{Loc}(\ell)$ for all $\ell^{\prime}\in B_{i_{j}}$. For the condition (iii), we have
$\begin{split}|B|=&\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|<\delta-1}|B_{i_{j}}|+\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|\geqslant\delta-1}|B_{i_{j}}|\\ =&T+\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|\geqslant\delta-1}(\delta-1)\\ \geqslant&T+(\delta-1)\left\lfloor\frac{|\tilde{B}|-T}{r+\delta-1}\right\rfloor\\ \geqslant&(\delta-1)\left\lfloor\frac{|\tilde{B}|}{r+\delta-1}\right\rfloor,\end{split}$ (12)
where
$T=\sum_{1\leqslant j\leqslant\tau-1\atop|\tilde{B}_{i_{j}}|<\delta-1}|B_{i_{j}}|$
and the first inequality holds by the fact that in each round the set $\tilde{B}$ reduces $|B_{i_{j}}|$ for the case $|\tilde{B}_{i_{j}}|<\delta-1$ and at most $|S_{i_{j}}|\leqslant r+\delta-1$ when $|\tilde{B}_{i_{j}}|\geqslant\delta-1$. This completes the proof. ∎
Now, with the results above, we proceed with the proof of the following improved lower bound on the access cost.
Theorem 2:
Let $\bm{\mathcal{C}}$ be a $(t,1)_{q}$ LRCC with parameters $(n_{I},k,*,*;n_{F},tk,r,\delta)_{a}$ where $\delta\geqslant 2$. Assume that the final code $\mathcal{C}_{F}$ is a linear code with $(r,\delta)_{a}$-locality and minimum distance $d$. Then, the write access cost is
$\rho_{w}\geqslant t\left(d+(t-1)k-1+\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1)\right)-(t-1)n_{F}.$ (13)
Denote $\Delta=n_{F}-2d-(t-1)k+2-(\lceil\frac{(t-1)k}{r}\rceil-1)(\delta-1)$. Then, the read access cost is
$\rho_{r}\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\\ t\left(k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),&\mbox{otherwise}.\end{cases}$
Moreover, if $d>n_{I}-k+1$, then the read access is at least $tk$.
Proof:
According to Lemma 4, the inequality (13) follows directly from $\rho_{w}=n_{F}-\sum_{i=1}^{t}|\mathcal{U}_{i,1}|$. Again, by Lemma 4, we have
$|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|\leqslant|\mathcal{U}_{i,1}|\leqslant n_{F}-d-(t-1)k+1-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1),$
which implies
$\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1\leqslant n_{F}-2d-(t-1)k+2-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1)\triangleq\Delta,$
where $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1$ is defined in Lemma 5. Thus, $\Delta_{i}\leqslant\Delta$ for every $i\in[t]$. According to Lemma 6, let $\tilde{B}=\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}$ then $|\tilde{B}|=\Delta_{i}+d-1$ by $\Delta_{i}=|\mathcal{U}_{i,1}\setminus\mathcal{R}_{i,1}|-d+1$. Then, we can find a $B$ satisfying condition (i)-(iii) and $|B|\geqslant(\delta-1)\lfloor\frac{\Delta_{i}}{r+\delta-1}\rfloor$. This is to say that
$\Delta_{i}-|B|$ $\leqslant\Delta_{i}-(\delta-1)\left\lfloor\frac{\Delta_{i}}{r+\delta-1}\right\rfloor$
$=(\delta-1)\left(\frac{\Delta_{i}}{\delta-1}-\left\lfloor\frac{\frac{\Delta_{i}}{\delta-1}}{\frac{r+\delta-1}{\delta-1}}\right\rfloor\right)$
$=(\delta-1)\left\lceil\frac{\left(\frac{r+\delta-1}{\delta-1}-1\right)\frac{\Delta_{i}}{\delta-1}}{\frac{r+\delta-1}{\delta-1}}\right\rceil$
$\leqslant(\delta-1)\left\lceil\frac{\left(\frac{r+\delta-1}{\delta-1}-1\right)\frac{\Delta}{\delta-1}}{\frac{r+\delta-1}{\delta-1}}\right\rceil$
$=\Delta-(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor.$
Then, by Lemma 5, we have
$|\mathcal{R}_{i,1}|\geqslant\begin{cases}k,&\mbox{if }\Delta\leqslant 0;\\ k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor,&\mbox{otherwise},\end{cases}$
for every $i\in[t]$. Therefore,
$\rho_{r}=\sum_{i=1}^{t}|\mathcal{R}_{i,1}|\geqslant\begin{cases}tk,&\mbox{if }\Delta\leqslant 0\\ t\left(k-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right),&\mbox{otherwise}.\end{cases}$
When $d>n_{I}-k+1$, the results follow directly from Lemma 5. ∎
Remark 3:
For the case $\delta=2$, the preceding bound is exactly the same as the one in Theorem 1 [23, Theorem IV.4]. While the case $\delta>2$, the new bound may outperform the one in Theorem 1. The reader may refer to Figure 1 as an example.
Figure 1: Compare the lower bounds on read access cost of [23, Theorem IV.4] and Theorem 2. We only consider the case $\Delta>0$ and $d\leqslant n_{I}-k+1$. Set the $t=3$ initial codes with parameters $[n_{I}=80,k=40]_{q}$, and the final code with parameters $[tk/\alpha,tk,d=15]_{q}$ and with $(r=10,\delta)_{a}$-locality where $\alpha$ is the code rate (the X-axis in the figure) of the final code. Note that the lower bound of [23, Theorem IV.4] is not relate to the $\delta$, therefore, the conditions for $\delta=3,4,5$ are the same blue circle in the figure.
Definition 4:
An LRCC is called access-optimal if and only if the read and write access cost both achieve the lower bound given by Theorem 2.
IV A General Construction of LRCC
In this section, we present a general construction for LRCCs with optimal access costs. The basic idea of our construction is to apply an original code to ensure a proper conversion from a set of codes, which are shortened versions of the original code. To begin, we prove the following simple lemma.
Lemma 7:
Let $g,r,\delta,\ell$ be positive integers with $g\geqslant\ell$. For an optimal $[g(r+\delta-1),\ell r,d]_{q}$-LRC $\mathcal{C}$ with $(r,\delta)_{a}$-locality, then $\mathcal{C}$ admits a parity check matrix $H$ of the form (1) where $A_{1},A_{2},\dots,A_{g}$ are $(\delta-1)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$, and $B_{1},B_{2},\dots,B_{g}$ are $((g-\ell)r)\times(r+\delta-1)$ matrices over $\mathbb{F}_{q}$. Then, for any set $\mathcal{P}=\{i_{1},\dots,i_{|\mathcal{P}|}\}\subseteq[g]$ with $|\mathcal{P}|>g-\ell$, the submatrix $H||_{\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $[|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
Proof:
The code $\mathcal{C}$ has the parity check matrix $H$ is directly from Lemma 3. Assume that a code $\tilde{\mathcal{C}}$ has the parity check matrix $H||_{\mathcal{P}}$. On the one hand, the minimum Hamming distance $d(\tilde{\mathcal{C}})$ of $\tilde{\mathcal{C}}$ satisfies $d(\tilde{\mathcal{C}})\geqslant d$, since any $d-1$ columns of $H||_{\mathcal{P}}$ are linear independent over $\mathbb{F}_{q}$. On the other hand, the dimension $k(\tilde{\mathcal{C}})$ of $\tilde{\mathcal{C}}$ satisfies
$k(\tilde{\mathcal{C}})$ $\geqslant|\mathcal{P}|(r+\delta-1)-(|\mathcal{P}|(\delta-1)+(g-\ell)r)$
$=(|\mathcal{P}|-g+\ell)r.$
Note that $\tilde{\mathcal{C}}$ has $(r,\delta)_{a}$-locality. According to Singleton-type bound (3), we have that
$d(\tilde{\mathcal{C}})$ $\leqslant|\mathcal{P}|(r+\delta-1)-(|\mathcal{P}|-g+\ell)r+1-(|\mathcal{P}|-g+\ell-1)(\delta-1)$
$=g(r+\delta-1)-\ell r+1-(\ell-1)(\delta-1)$
$=d,$
where the last equation holds by $\mathcal{C}$ is an optimal LRC. This is implies that $d(\tilde{\mathcal{C}})=d$, i.e., $\tilde{\mathcal{C}}$ is an optimal LRC with parameters $[|\mathcal{P}|(r+\delta-1),(|\mathcal{P}|-g+\ell)r,d]_{q}$. ∎
Theorem 3:
Let $g,\ell,r,\delta,t$ be positive integers with $t\mid\ell$, $\ell/t>g-\ell\geqslant 0$ and $\delta\geqslant 2$. Suppose $\mathcal{C}$ is an optimal $[g(r+\delta-1),\ell r,d]_{q}$-LRC with $(r,\delta)_{a}$-locality, then there exists an access-optimal LRCC with parameters
$((\ell/t+g-\ell)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$ (14)
for any non-negative integer $h\leqslant g-\ell$. Moreover, the initial and final codes are optimal LRCs.
Proof:
Let $H$ of the form (1) be the parity check matrix of $\mathcal{C}$. Denote $s=\ell/t$. Let $\mathcal{P}_{1},\dots,\mathcal{P}_{t},\mathcal{P}$ be the partition of $[g]$ such that $|\mathcal{P}_{i}|=s$ for all $i\in[t]$ and $|\mathcal{P}|=g-\ell$. Without loss of generality, assume $\mathcal{P}_{i}=\{(i-1)s+1,\dots,is\}$ for $i\in[t]$ and $\mathcal{P}=\{\ell+1,\dots,g\}$. According to Lemma 7 and $|\mathcal{P}_{i}\cup\mathcal{P}|=s+g-\ell$, the submatrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ defined by (2) is a parity check matrix for an optimal $[(s+g-\ell)(r+\delta-1),sr,d]_{q}$-LRC $\mathcal{C}_{i}$ with $(r,\delta)_{a}$-locality for $i\in[t]$. Let $\mathcal{C}_{i}$ be the initial code with the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$. The main part of the parity check matrix concatenating is the global parity check. Therefore, we first do a puncturing operation on the initial codes.
Without loss of generality, assume $i=1$. We divide $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ into blocks as follows:
$H||_{\mathcal{P}_{1}\cup\mathcal{P}}=\left(\begin{array}[]{ccc|ccc|ccc}A_{1}&\cdots&O&O&\cdots&O&O&\cdots&O\\ \vdots&\ddots&\vdots&O&\cdots&O&O&\cdots&O\\ O&\cdots&A_{s}&O&\cdots&O&O&\cdots&O\\ \hline\cr O&\cdots&O&A_{\ell+1}&\cdots&O&O&\cdots&O\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ O&\cdots&O&O&\cdots&A_{\ell+h}&O&\cdots&O\\ \hline\cr O&\cdots&O&O&\cdots&O&A_{\ell+h+1}&\cdots&O\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ O&\cdots&O&O&\cdots&O&O&\cdots&A_{g}\\ B_{1,1}&\cdots&B_{s,1}&B_{\ell+1,1}&\cdots&B_{\ell+h,1}&B_{\ell+h+1,1}&\cdots&B_{g,1}\\ \hline\cr B_{1,2}&\cdots&B_{2,2}&B_{\ell+1,2}&\cdots&B_{\ell+h,2}&B_{\ell+h+1,2}&\cdots&B_{g,2}\end{array}\right)=\left(\begin{array}[]{c|c|c}\bm{W}_{11}&\bm{O}&\bm{O}\\ \hline\cr\bm{O}&\bm{W}_{12}&\bm{O}\\ \hline\cr\bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \hline\cr\bm{N}_{11}&\bm{N}_{12}&\bm{N}_{13}\end{array}\right),$ (15)
where $B_{i,1}$ is consisting of the first $(g-\ell-h)r$ columns of $B_{i}$, and $B_{i,2}$ is consisting of the last $hr$ columns of $B_{i}$ for $i\in[g]$.
Note that $\bm{M}_{13}$ is a $(g-\ell-h)(r+\delta-1)\times(g-\ell-h)(r+\delta-1)$ matrix. Since $\mathcal{C}_{1}$ is optimal with the minimum Hamming distance $d=\delta+(g-\ell)(r+\delta-1)$, therefore,
[…]
There is an invertible square submatrix of $(\bm{M}_{13}^{\top},\bm{N}_{13}^{\top})^{\top}$ with order $(g-\ell-h)(r+\delta-1)$. Without loss of generality, let $\bm{M}_{13}$ be invertible. We derive a puncturing code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{1}$ by following elementary row operations on $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$:
$\begin{pmatrix}\bm{I}&\bm{O}&\bm{O}&\bm{O}\\ \bm{O}&\bm{I}&\bm{O}&\bm{O}\\ \bm{O}&\bm{O}&\bm{I}&\bm{O}\\ \bm{O}&\bm{O}&-\bm{N}_{13}\bm{M}_{13}^{-1}&\bm{I}\end{pmatrix}\cdot\begin{pmatrix}\bm{W}_{11}&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{12}&\bm{O}\\ \bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \bm{N}_{11}&\bm{N}_{12}&\bm{N}_{13}\end{pmatrix}=\begin{pmatrix}\bm{W}_{11}&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{12}&\bm{O}\\ \bm{M}_{11}&\bm{M}_{12}&\bm{M}_{13}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}&\bm{O}\end{pmatrix},$ (16)
where $\bm{I}$ is an identity matrix, $\bm{O}$ is a zero matrix.
Therefore, we have that the code $\mathcal{C}_{1}|_{[(s+h)(r+\delta-1)]}$ has parity check matrix
$\overline{H}_{1}=\begin{pmatrix}\bm{W}_{11}&\bm{O}\\ \bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}.$
Since every $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[t]$ has the same part $H||_{\mathcal{P}}$, thus the puncturing code $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ from $\mathcal{C}_{i}$ comes from the same elementary row operations, i.e., each $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ is left-multiplied by the first matrix in the left-hand side of (16). Then, $\mathcal{C}_{i}|_{[(s+h)(r+\delta-1)]}$ has the parity check matrix $\overline{H}_{i}$ for $i\in[t]$, i.e.,
$\overline{H}_{i}=\begin{pmatrix}\bm{W}_{i1}&\bm{O}\\ \bm{O}&\bm{W}_{12}\\ \bm{N}_{i1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{i1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix},$
which holds by $\bm{j}_{12}=\bm{j}_{22}=\cdots=\bm{j}_{t2}$ for $\bm{j}\in\{\bm{W},\bm{M},\bm{N}\}$, and $\bm{z}_{13}=\bm{z}_{23}=\cdots=\bm{z}_{t3}$ for $\bm{z}\in\{\bm{M},\bm{N}\}$.
To obtain the parity check matrix of the final code, we concatenate the parity check matrices $\overline{H}_{1},\dots,\overline{H}_{t}$ and derive the following matrix:
$H_{F}=\begin{pmatrix}\bm{W}_{11}&\bm{O}&\cdots&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{21}&\cdots&\bm{O}&\bm{O}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \bm{O}&\bm{O}&\cdots&\bm{W}_{t1}&\bm{O}\\ \bm{O}&\bm{O}&\cdots&\bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{21}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{21}&\cdots&\bm{N}_{t1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{t1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}.$ (17)
Now, we construct the final code $\mathcal{C}_{F}$. Choose any $t$ codewords $\bm{c}_{1}\in\mathcal{C}_{1},\dots,\bm{c}_{t}\in\mathcal{C}_{t}$, then we can calculate the new symbols as
$\bm{c}^{*}=\sum_{i=1}^{t}\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]},$ (18)
where for $i\in[h]$,
$A_{\ell+i}\cdot(\bm{c}^{*}|_{[i(r+\delta-1)]\setminus[(i-1)(r+\delta-1)]})^{\top}=\bm{0}$ (19)
and the code symbols $\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}$ corresponds to the read symbols for $\bm{c}_{i}$ for $1\leqslant i\leqslant t$. The final code $\mathcal{C}_{F}$ is defined as
$\mathcal{C}_{F}=\{(\bm{c}_{1}|_{[s(r+\delta-1)]},\dots,\bm{c}_{t}|_{[s(r+\delta-1)]},\bm{c}^{*}):\bm{c}_{1}\in\mathcal{C}_{1},\dots,\bm{c}_{t}\in\mathcal{C}_{t}\}.$
Let the code $\tilde{\mathcal{C}}$ be given by the parity check matrix $H_{F}$. We first show that $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$. On the one hand, according to (19) and the unchanged symbols satisfy the local parity check equations, for any codeword $\bm{c}\in\mathcal{C}_{F}$ and $i\in[\ell]$, we have
$(\underset{i-1\text{ times}}{\underbrace{O,\cdots,O}},A_{i},\underset{\ell+h-i\text{ times}}{\underbrace{O,\cdots,O}})\cdot\bm{c}^{\top}=\bm{0}.$
This implies that, for $i\in[t]$,
$(\underset{i-1\text{ times}}{\underbrace{\bm{O},\cdots,\bm{O}}},\bm{W}_{i1},\underset{t-i+1\text{ times}}{\underbrace{\bm{O},\cdots,\bm{O}}})\cdot\bm{c}^{\top}=\bm{0}.$ (20)
On the other hand,
$\begin{split}&\begin{pmatrix}\bm{O}&\cdots&\bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\cdots&\bm{N}_{t1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{t1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot\bm{c}^{\top}\\ &=\sum_{i=1}^{t}\begin{pmatrix}\bm{O}\\ \bm{N}_{i1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{i1}\end{pmatrix}\cdot(\bm{c}_{i}|_{[s(r+\delta-1)]})^{\top}+\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=-\sum_{i=1}^{t}\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}+\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot\left(\bm{c}^{*}-\sum_{i=1}^{t}\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}\right)^{\top}\\ &=\bm{0},\end{split}$ (21)
where the second equation holds by $\overline{H}_{i}\cdot\bm{c}_{i}|_{[(s+h)(r+\delta-1)]}^{\top}=\bm{0}$ for $i\in[t]$, and the last equation comes from (18). Combining (20) and (21), we have that $H_{F}\cdot\bm{c}^{\top}=\bm{0}$ for any $\bm{c}\in\mathcal{C}_{F}$. Therefore, $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$ and
$\ell r=\dim(\mathcal{C}_{F})\leqslant\dim(\tilde{\mathcal{C}}).$ (22)
Secondly, we claim that $\mathcal{C}_{F}$ is an optimal $[(\ell+h)(r+\delta-1),\ell r,d]_{q}$-LRC. By the fact $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$, it is sufficient to show that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$ and $\tilde{\mathcal{C}}$ is an optimal LRC. Note that $H_{F}$ is also the parity check matrix of the puncturing code $\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$ from $\mathcal{C}$ by (16) and (17). Thus, $\tilde{\mathcal{C}}=\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$. Recall that $\mathcal{C}$ is an optimal LRC, which means
$d=g(r+\delta-1)-\ell r+1-\left(\left\lceil\frac{\ell r}{r}\right\rceil-1\right)(\delta-1)=(g-\ell)(r+\delta-1)+\delta,$
by (3). Therefore, by Lemma 1, we have $\dim(\tilde{\mathcal{C}})=\ell r$ since $(g-\ell-h)(r+\delta-1)<d$, which together with (22) show that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$.
On the one hand, according to the puncturing operation on $\mathcal{C}$, the minimum Hamming distance of $\tilde{\mathcal{C}}$ satisfies
$d(\tilde{\mathcal{C}})\geqslant d-(g-\ell-h)(r+\delta-1)=\delta+h(r+\delta-1).$ (23)
On the other hand, $\tilde{\mathcal{C}}$ is an $[(\ell+h)(r+\delta-1),\ell r,d(\tilde{\mathcal{C}})]_{a}$-LRC with $(r,\delta)_{a}$-locality, by the Singleton-type bound (3), we have
$d(\tilde{\mathcal{C}})\leqslant(\ell+h)(r+\delta-1)-\ell r+1-\left(\left\lceil\frac{\ell r}{r}\right\rceil-1\right)(\delta-1)\\ =\delta+h(r+\delta-1).$ (24)
Combining (23) and (24), the code $\tilde{\mathcal{C}}=\mathcal{C}_{F}$ is an optimal LRC. Therefore, there exists an LRCC with parameters $((\ell/t+g-\ell)(r+\delta-1),\ell r/t,r,\delta;(\ell+h)(r+\delta-1),\ell r,r,\delta)_{a}$ as given in (14). The initial and final codes are all optimal LRCs.
Combining (18) and (19), we derive that the write access cost is $h(r+\delta-1)$ and the read access cost is $thr$. According to Theorem 2, we have that
$\rho_{w}$ $\geqslant t(\delta+h(r+\delta-1)+(t-1)sr-1+(s(t-1)-1)(\delta-1))$
$\quad-(t-1)(\ell+h)(r+\delta-1)$
$=h(r+\delta-1),$
and
$\Delta$ $=(\ell+h)(r+\delta-1)-2(\delta+h(r+\delta-1))$
$\quad-(t-1)sr+2-(s(t-1)-1)(\delta-1)$
$=(s-h-1)(r+\delta+1)+r$
$>0,$
where the last inequality holds by $s=\ell/t>g-\ell\geqslant h$. Therefore,
$\rho_{r}$ $\geqslant t\left(sr-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right)$
$=t(sr+h(r+\delta-1)-s(r+\delta-1)-\delta+1+(\delta-1)(s-h-1))$
$=thr.$
The read and write access costs both achieve the lower bound in Theorem 2. Thus, the LRCC above is access-optimal, which completes the proof. ∎
Remark 4:
The $t$ initial codes in Theorem 3 can be the same one code, when $h=|\mathcal{P}|=g-\ell$ and $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ for $i\in[t]$ where $D_{i}$ is an invertible diagonal matrix as
$D_{i}=\begin{pmatrix}D_{i1}&O_{s(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}.$
In this case, both $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ and $A_{i}\cdot H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ can be the parity check matrices of initial code $\mathcal{C}_{i}$, where
$A_{i}=\begin{pmatrix}D_{i1}&O_{s(\delta-1)\times h(\delta-1)}&O_{s(\delta-1)\times hr}\\ O_{h(\delta-1)\times s(\delta-1)}&I_{h(\delta-1)\times h(\delta-1)}&O_{h(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&O_{hr\times h(\delta-1)}&D_{i2}\end{pmatrix}.$
Therefore, for ever $1\leqslant i\leqslant t$,
$\overline{H}_{i}=H||_{\mathcal{P}_{i}\cup\mathcal{P}}=\begin{pmatrix}\bm{W}_{i1}&O\\ O&\bm{W}_{12}\\ \bm{N}_{i1}&\bm{N}_{12}\end{pmatrix}.$
And the conversion procedure function (18) to calculate the new symbols will becomes
$\bm{c}^{*}=\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}\end{pmatrix}^{-1}\sum_{i=1}^{t}\left(\begin{pmatrix}I_{h(\delta-1)\times h(\delta-1)}&O_{h(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}\right).$
Also the equation (20) and (21) hold in this case. Therefore, the new LRCC is access-optimal, and the initial and final codes are optimal LRCs.
V Explicitly Constructions of LRCCs
Recall that the proof of Theorem 3 gives the initial codes by means of the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ as (15) for $i\in[t]$, the final code by means of $H_{F}$ as (17), and the conversion procedure as (18). Also, the initial and final codes are all optimal LRCs, and the derived LRCC is access-optimal. In this section, we show two explicit constructions and it is sufficient to show the initial and final codes are all with super-linear length or are all maximally recoverable codes.
V-A Super-linear length
For MDS codes, the MDS conjecture implies that the length of MDS codes may not exceed $q+1$ ($q+2$ for $2\mid q$ and $k=3$), which means for MDS convertible codes the length of final codes may be upper bounded by $q+1$ ($q+2$ for $2\mid q$ and $k=3$). Analogously, Kong in [23] raises an interesting question about how small the field size can be to ensure the existence of LRCCs with optimal access costs.
The trivial answer for this question is the length of final codes for LRCCs can not exceed the optimal length of LRC referring the known bounds in [5, 24].
Lemma 8 ([5, Theorem 2]):
Let $n=w(r+\delta-1)$, $\delta\geqslant 2$, $k=ur+v$, and additionally, $r|k$ or $u\geqslant 2(r+1-v)$, where all parameters are integers. Assume there exists an optimal $[n,k,d]_{q}$ linear code $\mathcal{C}$ with all symbol $(r,\delta)$-locality, and define $t=\left\lfloor(d-1)/\delta\right\rfloor$. If $42t+1>$ 4, then
$n$$\leqslant\begin{cases}\frac{r+\delta-1}{r}\left(\frac{t-1}{2(q-1)}q^{\frac{2(w-u)r-2v-2}{t-1}}+1\right),&\text{ if }t\text{ is odd},\\ \frac{t(r+\delta-1)}{2r(q-1)}q^{\frac{2(w-u)r-2v}{t}},&\text{ if }t\text{ is even},\\ \end{cases}$
where $w-u$ can also be rewritten as $w-u=\lfloor(d-1+v)/(r+\delta-1)\rfloor$.
In this subsection, we propose an explicit construction, which implies that the above bound is also tight asymptotically for LRCCs for some special cases.
Base Code A ([24, Construction A]): Let $d\geqslant\delta+1$ and $n=m(r+\delta-1)$, where $d,\delta,r,m$ are positive integers. For $i\in[m]$, let $G_{i}=\{g_{i,1},g_{i,2},\dots,g_{i,r+\delta-1}\}$ be an $(r+\delta-1)$-subset of $\mathbb{F}_{q}$. Then, for each $i\in[m]$, we can construct a $(d-1)\times(r+\delta-1)$ Vandermonde matrix with generating set $G_{i}$ of the form $(U_{i}^{\top},V_{i}^{\top})^{\top}$, where
$U_{i}=\begin{pmatrix}1&1&\cdots&1\\ g_{i,1}&g_{i,2}&\cdots&g_{i,r+\delta-1}\\ \vdots&\vdots&\ddots&\vdots\\ g_{i,1}^{\delta-2}&g_{i,2}^{\delta-2}&\cdots&g_{i,r+\delta-1}^{\delta-2}\end{pmatrix},V_{i}=\begin{pmatrix}g_{i,1}^{\delta-1}&g_{i,2}^{\delta-1}&\cdots&g_{i,r+\delta-1}^{\delta-1}\\ g_{i,1}^{\delta}&g_{i,2}^{\delta}&\cdots&g_{i,r+\delta-1}^{\delta}\\ \vdots&\vdots&\ddots&\vdots\\ g_{i,1}^{d-2}&g_{i,2}^{d-2}&\cdots&g_{i,r+\delta-1}^{d-2}\end{pmatrix}.$
Let the code $\mathcal{C}$ have the following parity check matrix:
$H=\begin{pmatrix}U_{1}&O&\cdots&O\\ O&U_{2}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&U_{m}\\ V_{1}&V_{2}&\cdots&V_{m}\end{pmatrix}.$ (25)
The ensuing lemma shows sufficient conditions for Base Code A to be optimal.
Lemma 9 ([24, Theorem III.4]):
Let $r+\delta\geqslant d\geqslant 2\delta+1$. Suppose that for any subset $S\subseteq[m]$ with $2\leqslant|S|\leqslant\lfloor\frac{d-1}{\delta}\rfloor$, we have
$\left|\bigcup_{i\in S}G_{i}\right|\geqslant\left(r+\frac{\delta}{2}-1\right)|S|+\frac{\delta}{2},$ (26)
then any $d-1$ columns of $H$ given by (25) are linear independent. As a result, the base code $\mathcal{C}$ is an optimal $[n,k,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
Remark 5:
For more optimal locally repairable codes, the reader may refer to [18, 12, 37, 41, 1, 22, 28, 48, 5, 6, 7, 15, 39] and references therein.
Subsequent improved Singleton-type bound is employed to determine the dimension of the code.
Lemma 10 ([24, Remark II.4]):
When $d=r+\delta$ and $(r+\delta-1)\mid n$, the Singleton-type bound (3) can be improved as
$d\leqslant n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1).$
Therefore, we can derive the following corollary.
Corollary 1:
Let $m\geqslant v\geqslant 1$, and $i_{1}<i_{2}<\cdots<i_{v}$ where $i_{j}\in[m]$ for $j\in[v]$. Then the code $\mathcal{C}$ given by the parity check matrix $H||_{\{i_{1},i_{2},\dots,i_{v}\}}$ is an optimal $[v(r+\delta-1),(v-1)r,r+\delta]_{q}$-LRC with $(r,\delta)_{a}$-locality, where $H$ is the parity check matrix of the code $\tilde{\mathcal{C}}$ from Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (26).
Proof:
Since $G_{i},i\in[m]$ satisfies condition (26), then $G_{j},j\in{\{i_{1},i_{2},\dots,i_{v}\}}\subseteq[m]$ satisfies condition (26). Also, the $\overline{d}$ for $\mathcal{C}$ satisfies $\overline{d}=d=r+\delta\geqslant 2\delta+1$. According to Lemma 9 and Lemma 10, the dimension $k(\mathcal{C})$ of $\mathcal{C}$ with parameters $[v(r+\delta-1),k(\mathcal{C}),\overline{d}=r+\delta]_{q}$ satisfies
$r+\delta=v(r+\delta-1)-k(\mathcal{C})+1-\left\lceil\frac{k(\mathcal{C})}{r}\right\rceil(\delta-1)$
and
$k(\mathcal{C})\geqslant v(r+\delta-1)-(\delta-1)v-(\overline{d}-\delta)=(v-1)r.$
Therefore, $k(\mathcal{C})=(v-1)r$. ∎
Corollary 2:
Let $u$ be a positive integer. Let the code $\mathcal{C}$ be from Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (26). If $u\mid m-1$, then there exists an access-optimal LRCC with parameters
$((u+1)(r+\delta-1),ur,r,\delta;m(r+\delta-1),(m-1)r,r,\delta)_{a}.$
Moreover, the initial and final code are with super-linear length.
Proof:
Let $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-1}{u}]$, and $\mathcal{P}=\{m\}$. Then $\mathcal{C}_{i}$ given by the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[\frac{m-1}{u}]$ are the initial codes, where $H$ is given by (25). According to the proof of Theorem 3, the conversion procedure is (18) and the final code is $\mathcal{C}$. Therefore, there exists an access-optimal LRCC. According to Corollary 1, the initial codes are optimal $[(u+1)(r+\delta-1),ur,r+\delta]_{q}$-LRC with $(r,\delta)_{a}$-locality.
It is well-known that the base code $\mathcal{C}$ can be with super-linear length [24]. Therefore, the code length $m(r+\delta-1)$ satisfies that
$m(r+\delta-1)=\begin{cases}\Omega(q^{\delta}),&\mbox{if }d\geqslant 2\delta+1,\\ \Omega(q^{1+\frac{\delta}{2}-o(1)}),&\mbox{if }2\mid\delta,3\delta+1\leqslant d\leqslant 4\delta,\\ \Omega(q^{\frac{\delta+1}{2}}),&\mbox{if }2\nmid\delta,3\delta+1\leqslant d\leqslant 4\delta,\\ \Omega(q^{\frac{\delta}{2}}(q\log q)^{\frac{1}{\lfloor\frac{d-1}{\delta}\rfloor-1}}),&\mbox{if }2\mid\delta,d\geqslant 3\delta+1,\\ \Omega(q^{\frac{\delta}{2}}(q\log q)^{\frac{1}{2(\lfloor\frac{d-1}{\delta}\rfloor-1)}}),&\mbox{if }2\nmid\delta,2\mid\lfloor\frac{d-1}{\delta}\rfloor,d\geqslant 3\delta+1,\\ \Omega(q^{\frac{\delta}{2}}(q\log q)^{\frac{1}{2(\lfloor\frac{d-1}{\delta}\rfloor-2)}}),&\mbox{if }2\nmid\delta,2\nmid\lfloor\frac{d-1}{\delta}\rfloor,\lfloor\frac{d-1}{\delta}\rfloor>3,d\geqslant 3\delta+1.\end{cases}$
Since $(u+1)(r+\delta-1)=\frac{u+1}{m}\cdot m(r+\delta-1)$ and $\frac{u+1}{m}$ is a constant. Therefore, the initial codes can be with super-linear length. ∎
Remark 6:
For the case $2\delta+1=d$, $r=\delta+1$, the parameters in Lemma 8 can be figure out as $w-u=u+1-u=1$, $t=2$ and
$n\leqslant\frac{2\delta}{\delta+1}q^{\delta+1}/(q-1)=O(q^{\delta}).$
Thus, the code length $\Omega(q^{\delta})$ is asymptotically optimal with respect to the bound in Lemma 8.
Remark 7:
When arbitrary two of $\cup_{j\in\mathcal{P}_{i}}G_{j}$ for $i\in[t]$ are belong to different cosets of one multiplicative group of $\mathbb{F}_{q}^{*}$ in Corollary 2, i.e.,
$\beta_{i_{1}}^{-1}\cdot\bigcup_{j\in\mathcal{P}_{i_{1}}}G_{j}=\beta_{i_{2}}^{-1}\cdot\bigcup_{j\in\mathcal{P}_{i_{2}}}G_{j},$
where $i_{1}\neq i_{2},i_{1},i_{2}\in[t]$, and $\beta_{i}$ is a coset leader of the coset containing $\cup_{j\in\mathcal{P}_{i}}G_{j}$. Therefore, the $t$ initial codes can be the same one code by Remark 4. If the initial code is with super-linear length $n_{I}$, then the length of the final code is $\Theta_{t}(tn_{I})$ which can be super-linear.
V-B Maximally recoverable codes
In general, the local parity checks make it possible for LRCs to recover some predetermined erasure patterns whose size exceeds the minimum Hamming distance. However, it remains an open question whether LRCCs have a similar property. In this subsection, we apply explicit constructions to address this question. More specifically, we consider the case of maximally recoverable (MR) codes, which can recover from all information-theoretically recoverable erasure patterns.
Definition 5 ([13, 3]):
Let $\mathcal{C}$ be an $[n,k,d]_{q}$ code with $(r,\delta)_{a}$-locality, and define $\mathcal{S}=\{S_{i}:i\in[n]\}$, where $S_{i}$ is an $(r,\delta)$-repair set for coordinate $i$. The code $\mathcal{C}$ is said to be a maximally recoverable (MR) code if $\mathcal{S}$ is a partition of $[n]$, and for any $R_{i}\subseteq S_{i}$ such that $|S_{i}\setminus R_{i}|=\delta-1$, the punctured code $\mathcal{C}|_{\cup_{1\leqslant i\leqslant n}R_{i}}$ is an MDS code. Specially, if each $S_{i}\in\mathcal{S}$ is of size $|S_{i}|=r+\delta-1$, the code $\mathcal{C}$ is said to be an $(n,r,h,\delta,q)$-MR code, where $m=\frac{n}{r+\delta-1}$ and $h=mr-k$.
We introduce the linearized Reed-Solomon codes [30], where the main construction of MR codes in [3] is based on it.
Let $\mathbb{F}_{q}\subseteq\mathbb{F}_{q_{1}}$. For any $\alpha\in\mathbb{F}_{q_{1}}$ and $i\in\mathbb{N}$, the $\mathbb{F}_{q}$-linear operator $\mathcal{D}_{\alpha}^{i}:\mathbb{F}_{q_{1}}\to\mathbb{F}_{q_{1}}$ is defined by
$\mathcal{D}_{\alpha}^{i}(\beta)=\beta^{q^{i}}\prod_{j=0}^{i-1}\alpha^{q^{j}}.$
Let $\alpha\in\mathbb{F}_{q_{1}}$ and $\bm{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{h})\in\mathbb{F}_{q_{1}}^{h}$. For $i\in\mathbb{N}\cup\{0\}$ and $k,\ell\in\mathbb{N}$, where $\ell\leqslant h$, define the matrices
$D(\alpha^{i},\bm{\beta},k,\ell)=\begin{pmatrix}\beta_{1}&\beta_{2}&\cdots&\beta_{\ell}\\ \mathcal{D}_{\alpha^{i}}^{1}(\beta_{1})&\mathcal{D}_{\alpha^{i}}^{1}(\beta_{2})&\cdots&\mathcal{D}_{\alpha^{i}}^{1}(\beta_{\ell})\\ \vdots&\vdots&&\vdots\\ \mathcal{D}_{\alpha^{i}}^{k-1}(\beta_{1})&\mathcal{D}_{\alpha^{i}}^{k-1}(\beta_{2})&\cdots&\mathcal{D}_{\alpha^{i}}^{k-1}(\beta_{\ell})\end{pmatrix}\in\mathbb{F}_{q_{1}}^{k\times\ell}.$
In what follows, we introduce a construction for $(n,r,h,\delta,q)$-MR codes [3].
Base Code B([3, Construction A]): Let $\alpha_{1},\alpha_{2},\dots,\alpha_{r+\delta-1}$ be $r+\delta-1$ distinct elements in $\mathbb{F}_{q}^{*}$. Let
$P=\begin{pmatrix}1&1&\cdots&1\\ \alpha_{1}&\alpha_{2}&\cdots&\alpha_{r+\delta-1}\\ \vdots&\vdots&&\vdots\\ \alpha_{1}^{\delta-2}&\alpha_{2}^{\delta-2}&\cdots&\alpha_{r+\delta-1}^{\delta-2}\end{pmatrix}\in\mathbb{F}_{q}^{(\delta-1)\times(r+\delta-1)}.$
Let $\gamma_{1},\gamma_{2},\dots,\gamma_{h}\in\mathbb{F}_{q}^{h}$ form a basis of $\mathbb{F}_{q^{h}}$ over $\mathbb{F}_{q}$. Let $\bm{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{r+\delta-1})$ where $\beta_{i}=\sum_{j=1}^{h}\gamma_{j}\alpha_{i}^{\delta-2+j}$ for $i\in[r+\delta-1]$. For $m\in\mathbb{N}$, let $\mathcal{C}$ be the linear code with length $n$ over $\mathbb{F}_{q^{h}}$ given by the parity check matrix
$H=\begin{pmatrix}P&O&\cdots&O\\ O&P&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&P\\ D(\gamma^{0},\bm{\beta},h,r+\delta-1)&D(\gamma^{1},\bm{\beta},h,r+\delta-1)&\cdots&D(\gamma^{m-1},\bm{\beta},h,r+\delta-1)\end{pmatrix},$ (27)
where $\gamma\in\mathbb{F}_{q^{h}}$ is a primitive element.
Lemma 11 ([3, Theorem 2]):
Let $q\geqslant\max\{r+\delta,m+1\}$. Then the code $\mathcal{C}$ from Base Code B is an $(n=m(r+\delta-1),r,h,\delta,q^{h})$-MR code with minimum Hamming distance $d=(\lfloor\frac{h}{r}\rfloor+1)(\delta-1)+h+1$.
Therefore, we can derive the following corollary.
Corollary 3:
Let $q\geqslant\max\{r+\delta,m+1\}$, $m\geqslant v\geqslant\lceil\frac{h}{r}\rceil$, and $i_{1}<i_{2}<\cdots<i_{v}$ where $i_{j}\in[m]$ for $j\in[v]$. Then the code $\mathcal{C}$ given by the parity check matrix $H||_{\{i_{1},i_{2},\dots,i_{v}\}}$ is an $(n=v(r+\delta-1),r,h,\delta,q^{h})$-MR code, where $H$ is given by (27).
Corollary 4:
Let $u$ be a positive number and $q\geqslant\max\{r+\delta,m+1\}$. Let the code $\mathcal{C}$ be from Base Code B satisfying $r\mid h$, i.e., the dimension of $\mathcal{C}$ is a positive integer $k=(m-\frac{h}{r})r$. If $u\mid m-\frac{h}{r}$, then there exists an access-optimal LRCC with parameters
$((u+h/r)(r+\delta-1),ur,r,\delta;m(r+\delta-1),k,r,\delta)_{a}.$
Moreover, the initial and final codes are MR codes.
Proof:
Let $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-h/r}{u}]$, and $\mathcal{P}=\{m-h/r+1,\dots,m\}$. Then $\mathcal{C}_{i}$ given by the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[\frac{m-h/r}{u}]$ are the initial codes, where $H$ is given by (27). According to the proof of Theorem 3, the conversion procedure is (18) and the final code is $\mathcal{C}$ with the parity check matrix $H$. Therefore, there exists an access-optimal LRCC. According to Corollary 3, the initial codes are $((u+h/r)(r+\delta-1),r,h,\delta,q^{h})$-MR codes. ∎
Remark 8:
In Corollary 4, $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ where
$D_{i}=\begin{pmatrix}I_{u(\delta-1)\times u(\delta-1)}&O_{u(\delta-1)\times 1}&O_{u(\delta-1)\times 1}&\cdots&O_{u(\delta-1)\times 1}\\ O_{1\times u(\delta-1)}&1&0&\cdots&0\\ O_{1\times u(\delta-1)}&0&\prod_{j=0}^{1-1}(\gamma^{i-1})^{q^{j}}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ O_{1\times u(\delta-1)}&0&0&\cdots&\prod_{j=0}^{h-1}(\gamma^{i-1})^{q^{j}}\\ \end{pmatrix}.$
Therefore, the $t$ initial codes can be the same one code by Remark 4 when $h=r$.
VI Conclusion
This paper studies the code conversion of locally repairable codes in the merge regime. We improve the lower bound on access cost which is relevant to $(r,\delta)$. As a consequence, we give a general construction of access-optimal LRCCs. Then, based on the general construction, we show an explicit construction of access-optimal LRCCs such that the initial and final codes are with super-linear length. Also, the initial and final codes can be required with maximally recoverable property.